OpenStudy (anonymous):

10c-f=-13+cd solve for c. show work please

OpenStudy (anonymous):

i would start by isolating the variable c on one side

jhonyy9 (jhonyy9):

10c-cd=f-13 c(10-d)=f-13 c=(f-13)/(10-d)

OpenStudy (anonymous):

thx

jhonyy9 (jhonyy9):

yw